Find an equation of the line satisfying the given conditions
Through (0,5); m=-(2/3)
possible answers:
2x+3y=-15
2x-3y=15
3x+2y=-15
2x+3y=15

Answer by algebrapro18(249)

Find an equation of the line satisfying the given conditions Through (0,5);
m=-(2/3).

Well if we look closely we know that the y-intercept of the line is going to be 5 because when x = 0 y = 5. So we know that in slope intercept form b = 5. Now we just plug in what we know into slope intercept form.

y = mx+b
y = -2/3x +5

Now we need to get it into standard form.

y = -2/3x +5
3y = -2x + 15
2x + 3y = 15

Now you can check the answer.